1. (4 pts) Particles A and B start at the origin and move in the ry-plane. The velocity of Particle A is t = 7 - 31 and Parti 1. (4 pts) Particles A and B start at the origin and move in the xy-plane. The velocity of Particle A is−→u=−→i−3−→j and Particle B is moving in a direction perpendicular to Particle A. After 1 unit of time, the two particles become√26 units apart and Particle B is in the first quadrant, what is the velocity of Particle B?
Answer #1

Leti Vo = xi + Y3 are R VAIVO bence JAVno (1-35).(xi + vj) 20 x- 31-0 x=31 Now, V- Vauva Valz lv trol Vavavo (TAUB=0) (520)